how do u simplify 3 radical 8?
Answer by fractalier(6550)

Since 8 = 4 * 2, you can write
3sqrt(8) = 3 * sqrt(4) * sqrt(2)
3sqrt(8) = 3 * 2 * sqrt(2) = 6 * sqrt(2)